OpenStudy (anonymous):

simplify the expression: 1-sinx/cosx

OpenStudy (lgbasallote):

\[\huge (\frac{1-sinx}{cosx} )\mathbf{or} (1 - \frac{sinx}{cosx})\] i'd love for it to be the second one :P haha

sam (.sam.):

could be 1st one

OpenStudy (lgbasallote):

if it's the first one...just decompose... \[\huge \frac{1}{cosx} - \frac{sinx}{cosx}\] do you have an idea of the answer now?

OpenStudy (anonymous):

@lgbasallote yes i think so, u just separated them right?

OpenStudy (lgbasallote):

yeppp

OpenStudy (anonymous):

ok ^^ i got it